OpenStudy (anonymous):

Solve the quadratic equation x2 + 16x = -64 by factoring. A. {2, 4} B. 8 C. {-8, 8} D. -8

OpenStudy (anonymous):

add \(64\) and get \[(x^2+16x+64)=0\] this is a perfect square i.e. it is \[(x+8)^2=0\] solve that one

OpenStudy (anonymous):

this time there is only one solution, because if \(x+8=0\) then \(x=-8\) is the only choice